OpenStudy (anonymous):

i need to use two unit multiplers to make 250 inches to yards

OpenStudy (anonymous):

what is a unit multiplier

OpenStudy (sirm3d):

a unit multiplier converts one unit to another, like \[\frac{ 1ft }{ 12 inches }\]

OpenStudy (anonymous):

how would i convert inches to yards?

OpenStudy (sirm3d):

\[\left( \frac{ 1ft }{ 12 inch } \right)\left( \frac{ 1yd }{ 12ft } \right)\]

OpenStudy (sirm3d):

it is a two-step conversion, inch-to-foot then foot-to-yard. hence the two unit multipliers

OpenStudy (anonymous):

ooh thank you very much

OpenStudy (kropot72):

3 feet= 1 yard

OpenStudy (sirm3d):

@kropot72. thanks for the correction. i didn't see that. thanks again.

OpenStudy (kropot72):

You're welcome :)

hero (hero):

250 in = x yd 250 in = 3x ft 250 in = 36x in 250 = 36x 250/36 = x 6.944 = x